# Faster, Smaller, and Smarter: Task-Aware Expert Merging for Online MoE Inference

### v1

@@ Abstract
Sparse Mixture of Experts (SMoE) has become a preferred architecture for scaling Transformer capacity without increasing computational cost, as it activates only a small subset of experts for each input. However, deploying such an approach for online inference remains challenging due to the large size of a full SMoE model and the complexity of expert routing, especially in resource-constrained edge networks. Moreover, during the online inference, task information is often unavailable, making the task-level routing error-prone. In this work, we propose a novel tree-structured adaptive neural bandit router, Tanbr, to enable efficient and reliable online MoE inference. Instead of relying on explicit task tags, Tanbr estimates the task distribution over time from historical data and uses it to guide task-aware expert merging within a given pre-trained MoE. To handle the large continuous space of merging weights, Tanbr employs a binary tree to progressively partition the space and generate finer candidate weights. It then applies a neural bandit to learn the non-linear mapping from merging weight to model performance and decides optimal expert merging. We prove that Tanbr achieves a sublinear regret bound of $\mathcal{O}(\sqrt{T}\log(T))$ over $T$ rounds, despite operating over a continuous decision space, matching regret bounds compared to existing methods. Extensive experiments show that Tanbr reduces inference latency by at least $45\%$ and memory usage by up to $25\%$, while maintaining a high accuracy compared to many state-of-the-art methods.  [^1] Our code is provided at: https://anonymous.4open.science/r/Tanbr-MoE.
@@ I Introduction
Modern deep learning models, particularly large language models (LLMs), have shown remarkable performance in a variety of tasks [1, 2]. However, this advancement has come at the cost of increasing computational demands as model capacity has grown exponentially by approximately $10\times$ each year [3]. To address these scalability challenges, Sparse Mixture of Experts (SMoE) has emerged as a promising architecture. By routing inputs to a small subset of experts, SMoE models efficiently reduce training costs while improving accuracy for tasks such as language modeling [4], machine translation [5], and image classification [6]. Beyond training efficiency, the inference efficiency of MoE has received increasing attention due to the widespread deployment of large pre-trained models [7, 8]. To this end, task-level routing has emerged as a promising strategy. It assigns entire tasks to specific experts based on task tags, improving both training and inference efficiency [9]. Compared to token-level routing, this approach reduces the frequency of router activations and expert switching, thereby lowering inference latency. As shown in Tab. I (line 1 vs. 2 and 3 vs. 4), it achieves an efficiency gain of about $35\%$.
@@ III System Model
This section first outlines the preliminaries of a general MoE framework, and then we present the system model for effective online inference in an MoE system via expert merging.
Notation. We use lowercase letters to represent scalar values, bold letters for vectors and matrices. For any positive integer $I$, we use calligraphic letter $\mathcal{I}$ to denote the set $\{1,\ldots,I\}$.
@@ III-B Task-Aware Online Inference with MoE
Core Issue. To further reduce computational and memory overhead during inference, we also adopt expert merging strategy to serve inference tasks. The main challenge is to design and train a new router that can effectively leverage task features at each time slot to guide expert merging within the pre-trained MoE, while without requiring explicit task tags. Note that our router must dynamically adapt to evolving task distributions to ensure efficient and reliable online inference.
@@ III-C Problem Formulation
Challenges. i) Problem (1) can be reduced to the well-known Knapsack Problem [36], classifying it as an NP-hard problem. ii) The decision variables are both multi-dimensional and continuous, leading to high computational costs when searching for an optimal solution. iii) The reward acquisition is context-dependent and follows a non-linear mapping, which introduces additional complexities. iv) To preserve valuable information and utilize the problem’s linear components, the reward is multi-dimensional, which increases learning complexity and requires balancing multiple objectives.
@@ IV Algorithm Design
This section presents the tree-structured adaptive neural bandit router, Tanbr, which addresses the challenges of continuous and non-linear decision-making in router design.
@@ IV-A Main Idea
Given the challenges of Problem (1), we propose to utilize the MAB theory for learning and predicting unknown information. MAB-based methods are particularly well-suited, as they not only provide theoretical performance guarantees but also offer interpretability compared to conventional end-to-end learning approaches [37]. However, existing MAB methods are inadequate due to the presence of both non-linear parameters and a continuous decision space. Here, we introduce a novel framework, tree-structured adaptive neural bandit router, Tanbr, for fine-tuning and online inference within a pre-trained MoE. Tanbr includes the following steps:
@@ IV-C Design of Tanbr
18:  Update parameter: $\boldsymbol{\theta}_{t}=\boldsymbol{\theta}_{t-1}-\eta\nabla L(\boldsymbol{\theta}_{t-1})$;
@@
where $\boldsymbol{Z}$ is an auxiliary variable; $\gamma_{t}$ is a positive scaling factor and is defined as $\gamma_{t}=\sqrt{1+C_{1}w^{-1/6}\sqrt{\log w}L^{4}t^{7/6}\lambda^{-7/6}}\cdot(\upsilon\sqrt{\log(\det\boldsymbol{Z}_{t}/\det\lambda\boldsymbol{I})+C_{2}w^{-1/6}\sqrt{\log w}L^{4}t^{5/3}\lambda^{-1/6}-2\log\delta}$ $\sqrt{\lambda}S)+(\lambda+C_{3}tL)[(1-\eta dw\lambda)^{1/2}\sqrt{t/\lambda}+w^{-1/6}\sqrt{\log w}\cdot$ $L^{7/2}t^{5/3}\lambda^{-5/3}(1+\sqrt{t/\lambda})]$, where $\lambda$ denotes the regularization parameter, $S$ is the norm parameter, and $\eta$ represents the step size for updating the parameters of the neural network. The first term estimates the expected reward, the second reflects the granularity of the node’s region in the tree, and the third term quantifies the uncertainty in the reward prediction. Line 8 selects merging weight $\boldsymbol{x}_{h_{t},i_{t}}$ with the highest upper bound.
@@ IV-D Theoretical Analysis
$\left.+(\lambda+C_{3}TL)(1-\eta w\lambda)^{1/2}\sqrt{T/\lambda}+2\sqrt{\lambda S}\right]+T+1.$ (6)
Remark. By treating $\nu$, $\lambda$, and $\eta$ as constants and applying Lemma 1, the regret bound in Theorem 1 simplifies to $R_{T}=\mathcal{O}(\tilde{d}\sqrt{T}\log(T))$, matching the regret bounds in prior neural bandit studies [28, 29, 26, 27]. Compared to LinUCB [40], this bound includes an additional factor $\tilde{d}$, reflecting the complexity of neural representations. This extra cost is acceptable given the ability to capture more expressive and nonlinear context. A key contribution of Tanbr is maintaining these theoretical guarantees in an uncountable and continuous decision space. This is achieved through smoothness assumptions and adaptive partitioning techniques that ensure efficient exploration while avoiding exhaustive search. This result formally shows the theoretical suitability of Tanbr for long-term deployment in online MoE systems.
@@ V-A System Implementation
MoE Models and datasets. We evaluate our Tanbr using both i) encoder-decoder, and ii) decoder-only Transformer models to ensure a comprehensive evaluation. For both models, every FFN layer is replaced by an MoE layer containing $K=8$ experts with the same architecture. The encoder-decoder model is based on T5 [2], which includes $12$ FFN layers in both the encoder and decoder, and has about $1.0B$ parameters. The decoder-only model is based on BERT [1] and contains $12$ FFN layers with approximately $660M$ parameters. Both models are pre-trained using task routing [9, 33].
Datasets. We evaluate the models on $V=8$ common tasks from the GLUE benchmark, including natural language inference (i.e., MNLI [41], QNLI [42], RTE [43]), sentiment analysis (i.e., SST-2 [44]), sentence similarity (i.e., MRPC [45], QQP [46], STS-B [47]), and linguistic acceptability (i.e., CoLA [48]). Among them, CoLA is evaluated by the Matthews correlation coefficient, STS-B is evaluated by the average value of Pearson and Spearman correlation coefficients, and the remaining tasks are evaluated by accuracy. Each dataset is divided into training and validation sets in an $80/20$ split. Both the batch size and the maximum sequence length (for source and target) are set to $128$.
@@
Baselines. We compare Tanbr with the following methods: i) RL [49]: a reinforcement learning-based approach for expert selection in sparse MoE, assigning one expert per token. ii) Switch [4]: a neural network-based method that learns to select one expert per token for sparse MoE routing. iii) SMEAR [10]: a neural network-based token-level expert merging method. iv) NUCB [28]: a neural bandit method that randomly samples a set of 20 feasible merging weights and selects one by neural UCB at each step.v) Regmean [23]: a model merging method that computes weights using a closed-form solution to the merging problem. vi) Fisher [25]: a model merging method that leverages Fisher information to estimate parameter importance and performs weighted merging accordingly. vii) Average [50]: a model merging method that performs element-wise averaging of all parameters.
@@ V-B Results with Fine-tuning and Inference
Fine-tuning Phase. As mentioned in Section III, T5-based MoE with post-layer normalization requires fine-tuning when expert merging is applied due to its sensitivity to parameter shifts. Therefore, we fine-tune its normalization layers for $10,000$ steps using the training data, while simultaneously training a new router with different algorithms. The results are depicted in Fig. 2, showing the training loss curve, total training time, and memory usage. From the figure, we have some observations. First, Tanbr starts with higher loss since it begins with fewer candidate merging weights. However, as the partition tree refines over time, Tanbr converges much faster and achieves lower final loss than all baselines. In comparison, Nucb, which selects decisions from randomly generated merging weights, exhibits more fluctuation and slower convergence. Second, both Tanbr and Nucb (task-level routing) achieve the smallest training time, achieving approximately $40\%$ reduction compared to RL and Switch. Although Tanbr introduces a small amount of overhead due to maintaining the partition tree, it remains highly efficient. In contrast, Smear, which merges at the token level, results in a $20\times$ increase in computational cost because it cannot use lightweight adapters under the given pre-trained MoE architecture, making its overhead unavoidable. Third, Tanbr and Nucb consume about $10\%$ more memory than RL and Switch, as they need to maintain the loss gradients of multiple experts. Smear performs token-level merging, incurring the highest memory footprint, exceeding other methods by $179\%$ because it stores token-wise loss gradients for all experts.
@@ VI-B Proof of Theorem 1
$\left.+(\lambda+C_{3}TL)(1-\eta w\lambda)^{1/2}\sqrt{T/\lambda}+2\sqrt{\lambda S}\right]+T+1.$

### v2

# Ziyi Han1, Xutong Liu2, Ruiting Zhou3, Xiangxiang Dai1, John C.S. Lui1

@@ Abstract
Sparse Mixture of Experts (SMoE) has become a preferred architecture for scaling Transformer capacity without increasing computational cost, as it activates only a small subset of experts for each input. However, deploying such an approach for online inference remains challenging due to the large size of a full SMoE model and the complexity of expert routing, especially in resource-constrained edge networks. Moreover, during the online inference, task information is often unavailable, making the task-level routing error-prone. In this work, we propose a novel tree-structured adaptive neural bandit router, Tanbr, to enable efficient and reliable online MoE inference. Instead of relying on explicit task tags, Tanbr estimates the task distribution over time from historical data and uses it to guide task-aware expert merging within a given pre-trained MoE. To handle the large continuous space of merging weights, Tanbr employs a binary tree to progressively partition the space and generate finer candidate weights. It then applies a neural bandit to learn the non-linear mapping from merging weights to model performance and decides optimal expert merging. We prove that Tanbr achieves a sublinear regret bound of $\mathcal{O}(\sqrt{T}\log(T))$ over $T$ rounds, despite operating over a continuous decision space, matching regret bounds compared to existing methods. Extensive experiments show that Tanbr reduces inference latency by at least $45\%$ and memory usage by up to $25\%$, while maintaining a high accuracy compared to many state-of-the-art methods.
@@ I Introduction
Modern deep learning models, particularly large language models (LLMs), have shown remarkable performance in a variety of tasks [1, 2, 3]. However, this advancement has come at the cost of increasing computational demands as model capacity has grown exponentially by approximately $10\times$ each year [4]. To address these scalability challenges, Sparse Mixture of Experts (SMoE) has emerged as a promising architecture. By routing inputs to a small subset of experts, SMoE models efficiently reduce training costs while improving accuracy for tasks such as language modeling [5], machine translation [6], and image classification [7]. Beyond training efficiency, the inference efficiency of MoE has received increasing attention due to the widespread deployment of large pre-trained models [8, 9]. To this end, task-level routing has emerged as a promising strategy. It assigns entire tasks to specific experts based on task tags, improving both training and inference efficiency [10]. Compared to token-level routing, this approach reduces the frequency of router activations and expert switching, thereby lowering inference latency. As shown in Tab. I (line 1 vs. 2 and 3 vs. 4), it achieves an efficiency gain of about $35\%$.
@@ III-B Task-Aware Online Inference with MoE
Core Issue. To further reduce computational and memory overhead during inference, we also adopt expert merging strategy to serve inference tasks. The main challenge is to design and train a new router that can effectively leverage task features at each time slot to guide expert merging within the pre-trained MoE, without requiring explicit task tags. Note that our router must dynamically adapt to evolving task distributions to ensure efficient and reliable online inference.
@@ III-C Problem Formulation
Challenges. i) Problem (1) can be reduced to the well-known Knapsack Problem [37, 38], classifying it as an NP-hard problem. ii) The decision variables are both multi-dimensional and continuous, leading to high computational costs when searching for an optimal solution. iii) The reward acquisition is context-dependent and follows a non-linear mapping, which introduces additional complexities. iv) To preserve valuable information and utilize the problem’s linear components, the reward is multi-dimensional, which increases learning complexity and requires balancing multiple objectives.
@@ IV-A Main Idea
Given the challenges of Problem (1), reinforcement learning (RL) is less suitable, due to the absence of long-horizon state transitions [39]. We therefore adopt an MAB framework, which offers theoretical performance guarantees and greater interpretability compared to end-to-end learning methods[40, 41]. However, existing MAB methods are inadequate due to the presence of both non-linear parameters and a continuous decision space. Here, we introduce a novel framework, tree-structured adaptive neural bandit router, Tanbr, for fine-tuning and online inference within a pre-trained MoE. Tanbr includes the following steps:
@@ IV-C Design of Tanbr
where $\boldsymbol{Z}$ is an auxiliary variable; $\gamma_{t}$ is a positive scaling factor and is defined as $\gamma_{t}=\sqrt{1+C_{1}w^{-1/6}\sqrt{\log w}L^{4}t^{7/6}\lambda^{-7/6}}\cdot(\upsilon\sqrt{\log(\det\boldsymbol{Z}_{t}/\det\lambda\boldsymbol{I})+C_{2}w^{-1/6}\sqrt{\log w}L^{4}t^{5/3}\lambda^{-1/6}-2\log\delta}$ $\sqrt{\lambda}S)+(\lambda+C_{3}tL)[(1-\eta w\lambda)^{J/2}\sqrt{t/\lambda}+w^{-1/6}\sqrt{\log w}\cdot$ $L^{7/2}t^{5/3}\lambda^{-5/3}(1+\sqrt{t/\lambda})]$, where $\lambda$ denotes the regularization parameter, $S$ is the norm parameter, and $\eta$ represents the step size for updating the parameters of the neural network. The first term estimates the expected reward, the second reflects the granularity of the node’s region in the tree, and the third term quantifies the uncertainty in the reward prediction. Line 8 selects merging weight $\boldsymbol{x}_{h_{t},i_{t}}$ with the highest upper bound.
@@
18:  Update: $\boldsymbol{\theta}_{t}=\boldsymbol{\theta}_{t-1}-\eta\nabla L(\boldsymbol{\theta}_{t-1})$ with $J$ steps;
@@ IV-D Theoretical Analysis
$\left.+(\lambda+C_{3}TL)(1-\eta w\lambda)^{J/2}\sqrt{T/\lambda}+2\sqrt{\lambda S}\right]+T+1.$ (6)
Remark. Following Remark 4.7 in [29], by setting $J=\mathcal{O}(TL/\lambda)$ we have $(\lambda+C_{3}TL)(1-\eta w\lambda)^{J/2}\sqrt{T/\lambda}\leq\sqrt{\lambda S}$. By treating $\nu$, $\lambda$, and $\eta$ as constants and applying Lemma 1, the regret bound in Theorem 1 simplifies to $R_{T}=\mathcal{O}(\tilde{d}\sqrt{T}\log(T))$, matching the regret bounds in prior neural bandit studies [29, 30, 27, 28]. Compared to LinUCB [44], this bound includes an additional factor $\tilde{d}$, reflecting the complexity of neural representations. This extra cost is acceptable given the ability to capture more expressive and nonlinear context. A key contribution of Tanbr is maintaining these theoretical guarantees in a continuous decision space. This is achieved through smoothness assumptions and adaptive partitioning techniques that ensure efficient exploration while avoiding exhaustive search. This result formally shows the theoretical suitability of Tanbr for long-term deployment in online MoE systems.
@@ V-A System Implementation
MoE Models and datasets. We evaluate our Tanbr using both i) encoder-decoder, and ii) encoder-only Transformer models to ensure a comprehensive evaluation. For both models, every FFN layer is replaced by an MoE layer containing $K=8$ experts with the same architecture. The encoder-decoder model is based on T5 [2], which includes $12$ FFN layers in both the encoder and decoder, and has about $1.0B$ parameters. The decoder-only model is based on BERT [1] and contains $12$ FFN layers with approximately $660M$ parameters. Both models are pre-trained using task routing [10, 34].
Datasets. We evaluate the models on $V=8$ common tasks from the GLUE benchmark, including natural language inference (i.e., MNLI, QNLI, RTE), sentiment analysis (i.e., SST-2), sentence similarity (i.e., MRPC, QQP, STS-B), and linguistic acceptability (i.e., CoLA) [45, 46]. Among them, CoLA is evaluated by the Matthews correlation coefficient, STS-B is evaluated by the average value of Pearson and Spearman correlation coefficients, and the remaining tasks are evaluated by accuracy. Each dataset is divided into training and validation sets in an $80/20$ split. Both the batch size and the maximum sequence length (for source and target) are set to $128$.
@@
Baselines. We compare Tanbr with the following methods: i) RL [47]: a RL approach for expert selection in sparse MoE, assigning one expert per token. ii) Switch [5]: a neural network-based method that learns to select one expert per token for sparse MoE routing. iii) SMEAR [11]: a neural network-based token-level expert merging method. iv) NUCB [29]: a neural bandit method that randomly samples a set of 20 feasible merging weights and selects one by neural UCB at each step.v) Regmean [24]: a model merging method that computes weights using a closed-form solution to the merging problem. vi) Fisher [26]: a model merging method that leverages Fisher information to estimate parameter importance and performs weighted merging accordingly. vii) Average [48]: a model merging method that performs element-wise averaging of all parameters.
@@ V-B Results with Fine-tuning and Inference
Fine-tuning Phase. As mentioned in Section III, T5-based MoE with post-layer normalization requires fine-tuning when expert merging is applied due to its sensitivity to parameter shifts. Therefore, we fine-tune its normalization layers for $10,000$ steps using the training data, while simultaneously training a new router with different algorithms. The results are depicted in Fig. 2, showing the training loss curve, total training time, and memory usage. From the figure, we have some observations. First, Tanbr starts with a higher loss since it begins with fewer candidate merging weights. However, as the partition tree refines over time, Tanbr converges much faster and achieves lower final loss than all baselines. In comparison, Nucb, which selects decisions from randomly generated merging weights, exhibits more fluctuation and slower convergence. Second, both Tanbr and Nucb (task-level routing) achieve the smallest training time, achieving approximately $40\%$ reduction compared to RL and Switch. Although Tanbr introduces a small amount of overhead due to maintaining the partition tree, it remains highly efficient. In contrast, Smear, which merges at the token level, results in a $20\times$ increase in computational cost because it cannot use lightweight adapters under the given pre-trained MoE architecture, making its overhead unavoidable. Third, Tanbr and Nucb consume about $10\%$ more memory than RL and Switch, as they need to maintain the loss gradients of multiple experts. Smear performs token-level merging, incurring the highest memory footprint, exceeding other methods by $179\%$ because it stores token-wise loss gradients for all experts.
@@ VI-B Proof of Theorem 1
$\left.+(\lambda+C_{3}TL)(1-\eta w\lambda)^{J/2}\sqrt{T/\lambda}+2\sqrt{\lambda S}\right]+T+1.$